# Enhancing GANs with Contrastive Learning-Based Multistage Progressive Finetuning SNN and RL-Based External Optimization thanks: *GitHub repository: https://github.com/osamaa-mustafa/ImprovingGANs

### v2

@@ Abstract
Abstract..
Index Terms:
Deep Learning, Generative AI, Computer Visions, Generative Adversarial Networks, Contrastive Learning, Optimization, Reinforcement Learning, Cancer Research, Histopathology
I Abstract
The application of deep learning in cancer research, particularly in early diagnosis, case understanding, and treatment strategy design, emphasizes the need for high-quality data. Generative AI, especially Generative Adversarial Networks (GANs), has emerged as a leading solution to challenges like class imbalance, robust learning, and model training, while addressing issues stemming from patient privacy and the scarcity of real data. Despite their promise, GANs face several challenges, both inherent and specific to histopathology data. Inherent issues include training imbalance, mode collapse, linear learning from insufficient discriminator feedback, and hard boundary convergence due to stringent feedback. Histopathology data presents a unique challenge with its complex representation, high spatial resolution, and multiscale features. To address these challenges, we propose a framework consisting of two components. First, we introduce a contrastive learning-based Multistage Progressive Finetuning Siamese Neural Network (MFT-SNN) for assessing the similarity between histopathology patches. Second, we implement a Reinforcement Learning-based External Optimizer (RL-EO) within the GAN training loop, serving as a reward signal generator. The modified discriminator loss function incorporates a weighted reward, guiding the GAN to maximize this reward while minimizing loss. This approach offers an external optimization guide to the discriminator, preventing generator overfitting and ensuring smooth convergence. Our proposed solution has been benchmarked against state-of-the-art (SOTA) GANs and a Denoising Diffusion Probabilistic model, outperforming previous SOTA across various metrics, including FID score, KID score, Perceptual Path Length, and downstream classification tasks
II Introduction
Cancer especially breast cancer is a very common disease, in the UK and all around the world. Every year there are a large number of cases reported with the death rate eventually high. Cancer diagnosis and treatment is a complex and time-consuming process due to the multi-dimensional and non-static nature of cancer. Deep Learning has been utilized for tasks such as disease classification, tumor segmentation, and designing treatment strategies. This has helped the clinicians in early diagnosis and treatment of cancer. As in other domains, the speed of advancement towards Artificial General Intelligence (AGI) enabled with Multi-Modal Intelligence has been exponential, the speed has been observed to be comparatively slow in medical domain. Designing deep learning systems that are robust and transparent, it requires a large amount of data to train these systems. Data is a major concern in the medical domain due to patient privacy. Thus scarcity of data slows down the overall speed of Deep Learning advancement in this domain.
Generative Adversarial Networks (GANs) have been widely adopted for tasks such as synthetic data generation and style transfer [7]. Considering its success in other domains it has been extensively employed in the histopathology domain for tasks such as synthetic data generation and stain normalization. It has most widely been used for stain normalization and similar applications. An extensive literature review suggests, that researchers have done some remarkable work in the past in the field of generative histopathology. This includes the work of Runz et al [25] where they utilized cycle consistent generative adversarial networks (CycleGANs) to perform normalization of HE-stained histological images. In this work, Pegah Salehi et al. [26] employed a Pix2Pix network for stain-to-stain normalization in histopathology images. In another work, In this work [29], the authors have utilized a Generative Adversarial Network for stain normalization in federated learning based systems. GANs have been utilized for other auxiliary tasks such as for fibrosis detection and quantification in [22]. There are various other applications [3, 5, 16, 30, 18, 15, 23, 6, 32, 19, 28]. This is another interesting work where the authors Tibor Sloboda et al. [30] have performed editable stain transformation of histological images using unpaired GANs Most of the work is focused on style transfer and dealing with style and color in histopathology image data. Although GANs have been utilized extensively in this domain, there has been comparatively less work reported on synthetic image generation from noise. GANs have been previously used for tasks where the feature representation has been on a higher level but in histopathology images, due to high resolution and complex feature representation, it is comparatively difficult to train a model to generate histopathology images due to stability issues in GANs. A GAN learns to generate a synthetic image starting from random noise. There has been some work reported on GANs for image generation such as in this work [15] the authors Li et al. propose a multi-scale conditional GAN for image generation to be used in data augmentation. In another work on image generation [31], the authors propose HistoGAN for selective synthetic augmentation to improve downstream accuracy of classification.
Figure 1: An image illustrating a standard Architecture of a GAN.
However, GANs suffer from various issues although they have been the mainstream tool for data generation. Some of these issues are inherent to GANs which have been identified by the inventors themselves, while the other class of issues are those that have relevance to histopathology data such as complex feature representation. One of the major issues inherent to GAN that has been reported widely in the literature is the training imbalance and mode collapse. As in a GAN, the generator and discriminator train simultaneously in adversary to each other, they have their respective objective functions, the discriminator acts as a critic. As the generator generates a synthetic image out of random noise, the discriminator acts as a binary classifier and predicts whether the generated image is real or fake, and as the training progresses the generator tries to fool the discriminator. As the accuracy of the discriminator reaches at its highest, it does not let the discriminator learn further and it falls into local minima. This is the point where the generator starts to learn and generate noise. This is known as mode collapse. This has been referred to as ”The Helvetica Scenario” in the original GAN work by Goodfellow et el. [8]. In the same work the authors have also identified that the generator and discriminator should be well synchronized during training to avoid mode collapse. While it seems to be a simple solution but it is an active area of research as attaining this balance needs to be done by adding an auxiliary term which has also been identified as a future research direction in this work by the inventor Goodfellow et al. [8]. Therefore in our work, we propose a Reinforcement Learning powered External Optimizer (RL-EO) as an auxiliary term that guides the generator-discriminator adversary towards smooth convergence. Another issue that has been reported is that during training the discriminator is critic for the generator, based on feedback from the discriminator the gradients are calculated and a weight update is performed for the generator. But the discriminator in a standard GAN is a binary classifiers that does not provide sufficient information for proper convergence, as it just provides a binary feedback that whether the generated image is real or fake. This refers to if the discriminator does not learns well and starts to overfit the training data, it does not helps GAN in generating quality data which also leads to low generation quality. As been identified in the work by Goodfellow et al. [9], the model families we use are intrinsically flawed. Ease of optimization has come at the cost of models that are easily misled and many of the times the model is learning linear distribution of the input data and convex optimization-based training leads to easy convergence but linear nature. This issue relates to the inherent issue, that histopathology images are difficult to classify due to complex feature representation. Thus to solve this issue we introduce a representation learning-based Siamese Neural Network trained via contrastive loss to provide additional support to the discriminator. We modify the loss function of the discriminator to incorporate this. Some of the previous works have also experimented with integrating FID or Wasserstein Distance as a loss function in discriminator, and have identified that using these metrics actively in the training such as in loss function leads to the model learning linear distribution of the input data, thus overfitting to the input data.
Thus in our work, we propose solutions for major challenges reported in GANs. The major contributions are:
@@ II Introduction
The rest of the paper is structured as follows: Section III presents methodology, Section IV presents experimental protocol, followed by Section V & VI that presents evaluation, and results & discussion respectively. Section VII presents the conclusion.
III Methodology
This section details the methodology of the proposed framework. Our proposed solution strengthens the critic ability of the discriminator in the adversarial training by combining weakly supervised contrastive loss based SNN with GAN. As illustrated in Figure 4, we propose a novel framework in which we incorporate an external optimizer in the adversarial loop between the Generator and Discriminator in a standard GAN. The noise and training images are input to the pipeline, while the output from external optimizer (RL-EO) is passed as a weak reinforcement learning based reward signal to the discriminator. In the following sections, more details are provided explaining our proposed framework.
In order to provide external guidance to the discriminator to enable it to be more robust in acting as an adversary to the generator and to balance the generator discriminator training, we propose Mulltistage Finetuned Siamese Neural Network (MFT-SNN) as an external guide for convergence, acting as a reinforcement learning reward signaller for the discriminator. The proposed solution comprises of two parts where details for each part will be provided separately in the following sections.
  • Part-I: MFT-SNN: Multistage Progressive Finetuning based Siamese Neural Network.
  • Part-II: MFT-SNN as an RL-EO (Reinforcement Learning based External Optimizer) for the Generator-Discriminator training loop.
III-A MFT-SNN: Multistage Finetuning based Siamese Neural Network
@@
In this section, we provide further details on MFT-SNN’s training and configuration. Siamese Neural Networks [14] are a class of neural networks that serve the purpose of similarity measurement between a pair of input data. SNNs compute similarity by employing two identical weight-sharing networks. SNNs have been applied in tasks like signature forgery and document fraud detection, but limited studies report their use for similarity measurement in histopathology image data. The reason is that Histopathology images are very high resolution and multi-scale, thus they are first divided into patches. Thus the division of WSI into small patches increases the problems as the feature representation of a complete WSI is divided into patches. A challenge with histopathology data is that SNNs typically use convolutional networks to learn combined high- and low-level features, but in this case, the relevant information lies mostly in the low-level features.
@@
So in our proposed solution, we redefine the training strategy to handle all these challenges.
III-A1 Training Objective
SNN is provided pairs of images with labels 0 and 1 during training, indicating similarity and dissimilarity. Based on contrastive learning, the network is trained to minimize the distance between similar pairs and maximize the distance between dissimilar pairs.
III-A2 Feature Extractor
A Siamese Neural Network’s primary component is a feature extractor that is trained to extract feature embeddings from the pair of input images. For feature extraction, there are a range of available options from the class of convolutional architectures such as ResNet[18,34,40,101,152], InceptionNet, ResNeXt, and further variants. It is an option whether a pre-trained model is utilized or training is performed from scratch. Thus in our work, we employ a VGG-16 Autoencoder-based pre-trained architecture trained on ImageNet. Experimentation has been performed by employing various architectures and with/without pretraining. Experiments have also been performed comparing the performance of simple encoder-based architectures and encoder-decoder-based architectures. Results show that for this specific application, a pre-trained VGG-16 Autoencoder-based Siamese Neural outperforms its competitors based on performance. During the selection of feature extractor for SNN it is critical to consider the inference time complexity as later it gets integrated into the live training loop of GANs where two neural networks are being trained in adversary, the delay due to computational overhead affects the complete training process and creates a bottleneck.
III-A3 Proposed Training Strategy
In this section, details are provided regarding the proposed training strategy i.e multistage progressive fine-tuning based training of Siamese Neural Networks. The proposed strategy has two important components: Multistage and Progressive.
  • Multistage Training: This refers to the training process of the Siamese Neural Network (SNN) in two distinct stages. In the first stage, the SNN is trained on complete whole slide images, which are resized for computational efficiency without being divided into patches. After saving the weights from this stage, the same model is loaded to train on whole slide patches. In the second stage, the whole slide image is divided into patches, and the patch-level data is used to train the SNN.
  • Progressive Training: This refers to a two-stage training process in which the last layers of the pre-trained feature extractor are fine-tuned progressively. In the first stage, the last eight layers of the pre-trained feature extractor are fine-tuned using full-scale whole slide images. In the second stage, five of these eight layers are frozen, and the last three layers are fine-tuned again using patch-level data. Thus, the first stage involves fine-tuning all eight layers, while the second stage progressively fine-tunes the last three layers after freezing the first five.
@@ III-A3 Proposed Training Strategy
III-A4 Contrastive Learning
The Siamese Neural Network during training receives binary labels 0 and 1 for dissimilar and similar pairs of images. After training, when tested on test data with different levels of dissimilarity the model outputs a decreased similarity score with increased levels of dissimilarity. The model learned to quantify the similarity between images and shows a gradient of similarity scores that decrease with increasing dissimilarity, even though it was trained on binary labels. This indicates that the embedding space created by the Siamese network effectively represents the similarity continuum, which is a desirable outcome in contrastive learning.
III-B Mathematical Formulation for MFT-SNN Training
@@ 2. Stage 2 Optimization:
III-C Reinforcement Learning-Based External Optimizer (RL-EO)
This section details how the proposed MFT-SNN is integrated into GAN’s generator-discriminator training loop. It is a critical decision. In the proposed architecture, the Multistage Progressive finetuning-based Siamese Neural Network (MFT-SNN) is integrated as a Reinforcement Learning-Based External Optimizer (RL-EO) in the standard Generator-Discriminator training loop. In this work, the base GAN we selected is a standard GAN in its pure form, to ensure the transparency of performance increase due to the proposed modifications. To solve the variety of issues presented above leading to mode collapse and hard convergence, we integrate MFT-SNN as a valuable external guide for optimization. During training, the real and generated image is passed as an input pair to the MFT-SNN which computes the cosine similarity based on their extracted feature representations, the similarity score is a real-valued term between 0 and 1. In this work, we explored an ideal way to integrate the score into GAN. There are several ways it could be done such as passing it to the Generator module and passing it to the discriminator module. The similarity score is passed to the Discriminator.
@@ III-C Reinforcement Learning-Based External Optimizer (RL-EO)
The following subsection presents more details.
III-C1 RL-EO Output as Reward Signal to Discriminator
@@
III-C2 Modified Discriminator Loss Function
@@
where the similarity score is calculated using cosine similarity between the encoded representations of real and fake images.
@@ IV-A Dataset
Figure 7: Grad cam visualization - I
Figure 8: Grad cam visualization - II
IV-B Data Preparation
@@
IV-B1 Data for MFT-SNN
This part of the proposed solution is a Siamese Neural Network trained in two stages. For the first stage of training, the spatial resolution of WSIs is lowered down to 224x224x3 without dividing into patches. For the second stage of training, the WSIs are divided into patches. A total of 60,000 patches were selected for training in the second stage. The MFT-SNN requires images to be in pair format with binary labels assigned 0 or 1 corresponding to similar or dissimilar respectively. For training in the second stage, we prepared one similarity level and two dissimilarity levels as illustrated in Image 9. For each patch, we applied data augmentation including brightness, contrast, and noise. Different patches of same WSI and from different WSIs form two levels of dissimilarity.
Figure 9: Illustration of structural hierarchy for creation of similar and dissimilar pairs for SNN training.
IV-B2 Data for RL-EO Integrated GAN
This is the part of the proposed solution where the GAN is trained after the integration of Reinforcement Learning based External Optimizer (RL-EO). A total of 100,000 patches we selected for training the GAN.
IV-C Model Training
This section presents details regarding model training for the proposed solution including any architectures trained for benchmarking. This includes the train configuration set for training the models. All models are trained on a NVIDIA Tesla P100 16GB GPU hardware.
IV-C1 Training MFT-SNN
For MFT-SNN, the training is done in two stages as the proposed technique is a multistage progressive fine-tuning based SNN. For stage 1 of MFT-SNN training. a pretrained VGG-16 based Auto-encoder is fine-tuned on our dataset for 10 epochs. While the initial layers are freezed, only last 12 layers are allowed to be trainable. Further parameters are Batch Size: 32, Learning Rate for Reconstruction (Adagrad optimizer): 0.001, Learning Rate for Similarity (Adam optimizer): 0.0005. Two loss functions have been employed in this training: Contrastive loss for Siamese training and Mean Square Error as a reconstruction loss for the Autoencoder part.
The contrastive loss function is defined as:
$L_{\text{contrastive}}=\frac{1}{N}\sum_{i=1}^{N}\left[(1-y_{i})\cdot d_{i}^{2}+y_{i}\cdot\left(\max(0,m-d_{i})\right)^{2}\right]$
Where $N$ is the number of samples, $y_{i}$ is the label for the $i$-th pair (0 for similar pairs, 1 for dissimilar pairs), $d_{i}$ is the Euclidean distance between the pair of samples, and $m$ is the margin.
Breaking down the components:
  • $(1-y_{i})\cdot d_{i}^{2}$: This term penalizes similar pairs (when $y_{i}=0$) based on the squared Euclidean distance.
  • $y_{i}\cdot\left(\max(0,m-d_{i})\right)^{2}$: This term penalizes dissimilar pairs (when $y_{i}=1$) based on the squared margin minus the Euclidean distance, clamped to be non-negative.
For stage 2, we load the model fine-tuned in stage 1, freeze all the layers to preserve the training and only allow the last 3 layers to be trainable. The model is fine-tuned for 8 epochs with a batch size of 32. Learning rate used for the reconstruction optimizer (optimizer1) is set to 1e-3 (0.001). Learning rate used for the similarity optimizer (optimizer2) is set to 5e-4 (0.0005). Adagrad and Adam are optimizers respectively.
IV-C2 Training RL-EO Integrated GAN

### v3

@@ Abstract
Generative Adversarial Networks (GANs) have been at the forefront of image synthesis, especially in medical fields like histopathology, where they help address challenges such as data scarcity, patient privacy, and class imbalance. However, several inherent and domain-specific issues remain. For GANs, training instability, mode collapse, and insufficient feedback from binary classification can undermine performance. These challenges are particularly pronounced with high-resolution histopathology images due to their complex feature representation and high spatial detail.
In response to these challenges, this work proposes a novel framework integrating a contrastive learning-based Multistage Progressive Finetuning Siamese Neural Network (MFT-SNN) with a Reinforcement Learning-based External Optimizer (RL-EO). The MFT-SNN improves feature similarity extraction in histopathology data, while the RL-EO acts as a reward-based guide to balance GAN training, addressing mode collapse and enhancing output quality. The proposed approach is evaluated against state-of-the-art (SOTA) GAN models and demonstrates superior performance across multiple metrics.
1 Introduction
Breast cancer, with its high global incidence and mortality, poses significant diagnostic and treatment challenges. While Deep Learning supports early diagnosis through tasks like classification and tumor segmentation, progress in medical AI is hindered by data scarcity and privacy constraints, limiting effective training.
Generative Adversarial Networks (GANs) have been widely adopted for tasks such as synthetic data generation and style transfer [4]. Given its success in other domains, GANs have become widely used in histopathology for tasks such as synthetic data generation and stain normalization. Notable works include Runz et al.’s application of CycleGANs for color normalization [18], Salehi et al.’s use of Pix2Pix for stain-to-stain normalization [19], and GAN-based stain normalization within federated learning systems [21]. Additionally, GANs have shown promise in auxiliary histopathology tasks like fibrosis detection and quantification [16].
Most histopathology applications of GANs focus on style and color transfer to maintain data consistency; however, fewer studies address the direct generation of synthetic images from noise. Due to the high resolution and complex features of histopathology images, GANs face stability issues that make realistic image generation challenging. Recent advances, like Li et al.’s multi-scale conditional GAN for data augmentation [12] and HistoGAN for selective synthetic augmentation [22], have improved downstream classification, but challenges remain.
Despite the popularity of GANs in data generation, significant challenges persist, especially with high-complexity histopathology data. Training imbalance and mode collapse occur when adversarial dynamics fail to balance the generator and discriminator, leading to noise generation—a phenomenon referred to by Goodfellow et al. as the “Helvetica Scenario” [5]. Our approach addresses these issues with a Reinforcement Learning-based External Optimizer (RL-EO) to guide smoother convergence.
Another major limitation is the discriminator’s binary feedback, which can lead to overfitting and restricts GANs to linear data distributions unsuitable for complex histopathology features [6]. To improve the discriminator’s robustness, we incorporate a representation learning-based Siamese Neural Network (SNN), trained with contrastive loss, to provide refined feedback on generated image quality. While researchers have previously tried incorporating FID and Wasserstein Distance as integrated loss metrics, they lead to clear overfitting thus model learning linear distribution of input data, highlighting the need for our approach. The major contributions are:
@@ 1 Introduction
The rest of the paper is structured as follows: Section 2 presents methodology, Section 3 presents experimental protocol, followed by Section 4 & 5 that presents evaluation, and results & discussion respectively. Section 6 presents the conclusion.
2 Methodology
The proposed solution strengthens the critic ability of the discriminator. As illustrated in Figure 3, an external optimizer is integrated in the adversarial loop between the Generator and Discriminator in a standard GAN. The output from the external optimizer (RL-EO) is passed as a weak reinforcement learning-based reward signal to the discriminator.
We propose a Mulltistage Progressive Finetuned Siamese Neural Network (MFT-SNN) as an external guide for convergence, acting as a reinforcement learning reward signaller for the discriminator. The proposed solution comprises of two parts where details for each part will be provided separately in the following sections.
2.1 MFT-SNN: Multistage Finetuning based Siamese Neural Network
@@
Siamese Neural Networks (SNNs) [11] face challenges in histopathology due to high-resolution images being segmented into patches, which complicates the capture of relevant low-level features.
@@
So in our proposed solution, we redefine the training strategy to handle these challenges.
2.1.1 Feature Extractor and Training Objective
The SNN is trained on image pairs labeled 0 (dissimilar) and 1 (similar) to minimize the distance for similar pairs and maximize it for dissimilar pairs based on contrastive learning. In our work, we focus on using a VGG-16 Autoencoder pre-trained on ImageNet as the feature extractor in a Siamese Neural Network for generating embeddings from input image pairs. While various convolutional architectures like ResNet, InceptionNet, and ResNeXt can be utilized, our experiments demonstrate that the VGG-16 Autoencoder outperforms these alternatives for this application. We also emphasize the importance of inference time complexity in feature extractor selection, as computational overhead can create bottlenecks during live training of GANs.
2.1.2 Proposed Training Strategy
The proposed training strategy has two important components: Multistage and Progressive.
  • Multistage Training: This refers to the training process of the Siamese Neural Network (SNN) in two distinct stages. First, the SNN is trained on complete whole slide images, which are resized for computational efficiency without being divided into patches. In the second stage, the whole slide image is divided into patches, and the patch-level data is used to train the SNN.
  • Progressive Training: This refers to a two-stage training process in which the last layers are unfrozen for fine-tuning progressively. In the first stage, the last eight layers of the pre-trained feature extractor are fine-tuned using full-scale whole slide images. In the second stage, five of these eight layers are frozen, and the last three layers are fine-tuned again using patch-level data. Thus, the first stage involves fine-tuning all eight layers, while the second stage progressively fine-tunes the last three layers after freezing the first five.
@@ 2.1.2 Proposed Training Strategy
2.2 Mathematical Formulation for MFT-SNN Training
@@ 2. Stage 2 Optimization:
2.3 Reinforcement Learning-Based External Optimizer (RL-EO)
This section details how the proposed MFT-SNN is integrated into GAN’s generator-discriminator training loop. It is a critical decision. In the proposed architecture, the Multistage Progressive finetuning-based Siamese Neural Network (MFT-SNN) is integrated as a Reinforcement Learning-Based External Optimizer (RL-EO) in the standard Generator-Discriminator training loop. In this work, the base GAN we selected is a standard GAN in its pure form, to ensure the transparency of performance increase due to the proposed modifications. To solve the variety of issues presented above leading to mode collapse and hard convergence, we integrate MFT-SNN as a valuable external guide for optimization. During training, the real and generated image is passed as an input pair to the MFT-SNN which computes the cosine similarity based on their extracted feature representations, the similarity score is a real-valued term between 0 and 1. In this work, we explore an ideal way to integrate the score into GAN. There are several ways it could be done such as passing it to the Generator module and passing it to the discriminator module. The similarity score is passed to the Discriminator.
@@ 2.3 Reinforcement Learning-Based External Optimizer (RL-EO)
2.3.1 RL-EO Output as Reward Signal to Discriminator
@@
2.3.2 Modified Discriminator Loss Function
@@ 3.1 Dataset
Figure 6: Sample WSI-I
Figure 7: Sample WSI-II
3.2 Data Preparation
@@
3.2.1 Data for MFT-SNN and Modified GAN
For the first stage, WSIs are resized to 224x224x3 without patch division. In the second stage, 60,000 patches are selected for training, requiring paired images with binary labels (0 for dissimilar, 1 for similar). We established one similarity level and two dissimilarity levels, applying data augmentation (brightness, contrast, and noise) to create variations among patches from the same and different WSIs. A total of 100,000 patches were selected for training the RL-EO integrated GAN.
3.3 Model Training
This section presents details regarding model training for the proposed solution including any architectures trained for benchmarking. This includes the train configuration set for training the models. All models are trained on a NVIDIA Tesla P100 16GB GPU.
3.3.1 Training MFT-SNN
For MFT-SNN, the training is done in two stages as the proposed technique is a multistage progressive fine-tuning based SNN. For stage 1 of MFT-SNN training. a pretrained VGG-16 based Auto-encoder is fine-tuned on our dataset for 10 epochs. While the initial layers are freezed, only last 12 layers are allowed to be trainable. Further parameters are Batch Size: 32, Learning Rate for Reconstruction (Adagrad optimizer): 0.001, Learning Rate for Similarity (Adam optimizer): 0.0005. Two loss functions have been employed in this training: Contrastive loss for Siamese training and Mean Square Error as a reconstruction loss for the Autoencoder part. For stage 2, we load the model fine-tuned in stage 1, freeze all the layers to preserve the training and only allow the last 3 layers to be trainable. The model is fine-tuned for 8 epochs with a batch size of 32. Learning rate used for the reconstruction optimizer (optimizer1) is set to 1e-3 (0.001). Learning rate used for the similarity optimizer (optimizer2) is set to 5e-4 (0.0005). Adagrad and Adam are optimizers respectively.
3.3.2 Training RL-EO Integrated GAN